# Scizor: A Self-Supervised Approach to Data Curation for Large-Scale Imitation Learning

### v1

@@ 4.1 Experimental Setup
  • DemoInf [6]: A trajectory-level method that estimates mutual information between states and actions for each trajectory as a quality score and removes low-quality trajectories with insufficient contribution. We apply it to the RoboMimic and the Sirius-Fleet dataset, following the same dataset settings as in the original paper. We delete the same percentage of data as other methods.
@@ 4.2 Experimental Results
Figure 3: Performance comparison across different datasets. We use the unified threshold for Scizor and report success rates on $4$ datasets. We found that Scizor achieves the strongest performance and outperforms the baselines.
@@
RQ3: What design components contribute most to Scizor? We first ablate the suboptimal data removal and the deduplication in Table 1. We run the experiments only removing suboptimal data and duplicated data, and remove the same amount of data in each dataset as Scizor. We find that both suboptimal removal and deduplication individually lead to improvements over the baseline, but neither alone is sufficient to match the full performance of Scizor. Suboptimal removal is generally more effective than deduplication, but combining both components leads to the largest gains across all datasets.
@@
Figure 4: Breakdown of suboptimal types classified by Scizor. The three dominant failure modes predicted by Scizor’s suboptimal classifier are Slow Motion, Manipulation Failure, and Pause, showing Scizor removes semantically meaningful transitions.
RQ4: What types of low-quality samples can Scizor identify and curate? To qualitatively visualize the suboptimal data identified, we examine the low-quality data classified and investigate the types of low-quality data they represent. From each of the RoboMimic and Sirius-Fleet datasets, we randomly select $100$ demonstrations flagged with at least one suboptimal segment. We then manually visualize and classify every suboptimal segment across these trials to generate the pie chart. Figure 4 illustrates the distribution of suboptimal transitions identified by Scizor. Manipulation Failure refers to errors during grasping — e.g., failed grasps, accidental drop of objects. Pause denotes transitions with very little or no movement. Stuck at Collision describes cases where the gripper or held object collides, leading to a halt. Slow captures motion that proceeds noticeably below the normal task speed. Move Back and Forth indicates aimless motions that do not contribute to task progress. False Positive labels misclassified or ambiguous transitions. The most significant fractions are Slow Motion, Manipulation Failure, and Pause, indicating that the task-progress classifier identifies semantically meaningful errors rather than spurious noise. Appendix B.4 visualizes representative examples.
@@ 6 Limitation
Curation Threshold: The deletion threshold in Scizor is currently determined empirically. Future work could derive a theoretical framework to identify this threshold more systematically.
@@
Linear Task Progress Assumption: Scizor assumes linear task progression without slow or repetitive behaviors. However, real-world tasks, like stirring food repeatedly or waiting for it to cook, often involve such behaviors. Future work could adapt the method to better handle these behaviors.
@@ Appendix A Method details
A.1 Determining the Unified Threshold
@@ B.2 Training and Evaluation Details
Table 4 provides the detailed hyperparameters and model architectures used in our experiments.
B.3 Evaluation Task Details for each Dataset
@@
B.4 Extra results and Visualization
Table 6, 7, 8 presents the detailed results of different methods on each task of the reported datasets.
@@ B.4 Extra results and Visualization
B.5 Findings during Evaluation

### v2

@@ 4.1 Experimental Setup
  • DemoInf [6]: A trajectory-level method that estimates mutual information between states and actions for each trajectory as a quality score and removes low-quality trajectories.
@@ 4.2 Experimental Results
Figure 3: Performance comparison across different datasets. We use the unified threshold for Scizor and report success rates on 44 datasets. We found that Scizor achieves the strongest performance and outperforms the baselines.
@@
RQ3: What design components contribute most to Scizor? We first ablate suboptimal data removal and deduplication in Table 1. We run experiments only removing suboptimal data or duplicated data, and remove the same amount of data in each dataset as Scizor. We find both suboptimal removal and deduplication individually lead to improvements over the baseline, but neither alone is sufficient to match the full performance of Scizor. Suboptimal removal is generally more effective than deduplication, but combining both components leads to the largest gains across all datasets.
@@
Figure 4: Breakdown of suboptimal types classified by Scizor. The three dominant failure modes predicted by Scizor’s suboptimal classifier are Laggy Motion, Manipulation Failure, and Pause, showing Scizor removes semantically meaningful transitions.
RQ4: What types of low-quality samples can Scizor identify and curate? To qualitatively visualize the suboptimal data identified, we examine the predicted low-quality data and investigate the types of low-quality data they represent. From each of the RoboMimic and Sirius-Fleet datasets, we randomly select $100$ demonstrations flagged with at least one suboptimal segment. We then manually visualize and classify every suboptimal segment across these trials to generate the pie chart. Figure 4 illustrates the distribution of suboptimal transitions identified by Scizor. Manipulation Failure refers to errors during grasping—for example, failed grasps, accidental drop of objects. Pause denotes transitions with no movement. Stuck at Collision describes cases where the gripper or held object collides, leading to a halt. Lagging captures motion that proceeds noticeably below the normal speed for the same context. This doesn’t contain motions that are intentionally slow, careful, and deliberate. Move Back and Forth indicates aimless motions that don’t contribute to task progress. False Positive labels misclassified or ambiguous transitions. The most significant fractions are Laggy, Manipulation Failure, and Stalled, indicating that the task-progress classifier identifies semantically meaningful errors rather than spurious noise. Appendix B.6 visualizes representative examples.
@@ 6 Limitation
Linear Task Progress Assumption: Scizor assumes linear task progression without pausing or repetitive behaviors. However, real-world tasks, like stirring food repeatedly or waiting for it to cook, often involve such behaviors. Future work could adapt the method to better handle these behaviors, e.g., longer history input, hierarchical progress modeling that can capture sub-tasks, multi-timescale progress models that can capture both long-term progress and short-term progress, etc.
@@ Appendix A Method details
A.1 Determining the Single Threshold Across Datasets
When curating data, we remove all samples with a score above either $\epsilon_{s}$ or $\epsilon_{d}$. To generalize Scizor to different datasets, we suggest to find a single threshold. We conduct a hyperparameter search on two simulation datasets, RoboMimic and OXEMagic and find that $\epsilon{s}=0.58$ and $\epsilon_{d}=0.99$ yield the best performance on both datasets. This single threshold was then directly applied to the real-world setting (Sirius-Fleet).
@@ A.4 Fixed Time Duration
A.5 Task Progress Predicting Details
Rather than regressing a real-valued progress estimate, we cast progress prediction as classification over discrete temporal bins, which is empirically more robust [16]. We discretize the temporal gap into $B$ bins, where each bin is a time interval in seconds. To predict task progress for a sub-trajectory $S_{i,i+T}$, we train a task progress classifier to classify the bin corresponding to the time $T$ between the start and end states. Empirically, we set $B=5$, sample each sub-trajectory as $2$ seconds, and bins to be $[0,0.5),[0.5,1.0),[1.0,2.0),[2.0,5.0)$, and $5.0,+\infty)$. For the weighted combination coefficient between local suboptimality score and the mean score across all time steps, we simply choose 0.5.
A.6 Removing Suboptimal Samples Details
Note that each sample preserves an observation history and an action sequence for algorithms that are history-dependent (e.g., BC-Transformer) and that utilize action chunking. We do the chunking before any removal to avoid temporal inconsistencies.
A.7 Deduplication Details
We use a chunk size of N=8 for deduplication, as it’s commonly supported by most of the video encoder.
@@ B.2 Training and Evaluation Details
For each setting, the same dataset is used throughout SCIZOR suboptimal classifier training, deduplication, and subsequent policy training. For example, in the OXE setting, we first train SCIZOR’s task progress predictor on the full OXE dataset and use it to identify suboptimal transitions within that dataset. After filtering out low-quality segments, we apply de-duplication to the remaining data. The final policy is then trained on this curated dataset. Our pipeline is entirely self-supervised, requiring no explicit data quality labels for training.
@@
Table 4 provides the detailed hyperparameter and model architectures used in our experiments.
B.3 Other Simpler Baselines
We further compare Scizor against two simple baselines. To show the necessity of progress estimation, we design a velocity filter: we compute the end-effector velocity as the L2 norm of the delta end effector state and apply a moving-window filter to remove the slowest k% of transitions. To evaluate whether zero-shot VLMs can serve as effective demonstration raters, we prompt GPT-4.1-small to categorize each demonstration into three ranks and discard the lowest k%. Results in Table 5 show that only Scizor improves over the No Curation baseline. The moving window velocity filter catches only slow movements; once those are deleted, further pruning harms performance. The VLM baseline helps on the “Can” task but hurts on “Square”, indicating its limited spatial understanding for precise, fine-grained assembly.
Velocity Filter Zero-Shot VLM No Curation Scizor
Can (%) $75.7\pm 1.3$ $80.4\pm 4.0$ $79.0\pm 1.7$ $\mathbf{87.3\pm 0.7}$
Square (%) $30.1\pm 3.2$ $32.1\pm 2.4$ $34.8\pm 5.1$ $\mathbf{37.2\pm 2.5}$
Table 5: Comparison of methods on Robomimic.
B.4 Additional Research Questions
RQ5: What are the false negatives and false positives? We analyze the suboptimal cases that were incorrectly classified by Scizor, labeled as false positive in Figure 4. Among the $5\%$ of false positive segments predicted in Sirius-Fleet, $3\%$ occur near the end of the demonstrations. These segments often include repeated padded frames, since the task has already been completed. The remaining 2% are short segments right before a suboptimal pause, which would also be predicted as suboptimal segments. To conduct a controlled error analysis, we created a human-annotated dataset of robot failures from the Sirius-Fleet dataset. 13/25 of failures are successfully predicted. The false negatives can be categorized into two types: (1) The robot failed to grasp at first, but it recovered quickly and didn’t cause a big progress lag. This is because our curation threshold focuses on deleting the most detrimental suboptimality. (2) The self-supervised model failed to detect because most of the data contains this failure.
RQ6: How well can Scizor preserve high quality data? We investigate whether SCIZOR removes high-quality data during the curation process through the following analyses. 1) Keeping Expert Demonstrations. We randomly picked 50 perfect expert demonstrations from 5 tasks from Sirius-Fleet, each containing 10 demonstrations. Surprisingly,none of these demonstrations have SCIZOR-predicted suboptimal segments, indicating that SCIZOR is actually considering them as high-quality samples and won’t be curated. 2) Keeping Recovery Behaviors. We also checked the grasping failures mentioned in the previous paragraph. Specifically, 54% of recovery behaviors are fully maintained, and 31% are partially maintained. Only 15% are deleted, and the remaining episodes directly start with the next rasping behavior. Most of the recovering behaviors are kept because they contribute to task progress.
B.5 Evaluation Task Details for each Dataset
@@
B.6 Extra results and Visualization
Table 7, 8, 9 presents the detailed results of different methods on each task of the reported datasets.
@@ B.6 Extra results and Visualization
B.7 Findings during Evaluation